# Requirements Coverage-Guided Minimization for Natural Language Test Cases

### v1

@@ 4.5. Evaluation Metrics
(11) $Coverage=\frac{|R^{\prime}|}{|R|}$
where $R^{\prime}$ represents the set of requirements covered by the minimized test suite $T^{\prime}$, and $R$ denotes the original set of requirements. The coverage value ranges from 0 to 1, where 0 indicates that the minimized test suite covers no requirements, while 1 signifies full coverage of all requirements.
@@ 4.6. Experiment Environment
All experiments were conducted on a computer provided by our industry partner, running Windows 10, equipped with an Intel Core i7-11850H CPU at 2.5 GHz and 32 GB of RAM. Due to data privacy considerations, we were restricted to performing our experiments solely on this laptop, which limited the choice of text representation techniques available to us.
@@ 5.1. Performance of RTM under different configurations (RQ1)
Approach. To address RQ1, we systematically evaluated the performance of RTM under various configurations. Specifically, under a mid-range minimization budget ($50\%$), we ran RTM with three different preprocessing strategies (as described in Section 3.1), and applied both word-level and sentence-level embedding techniques to transform test cases into vector representations (Section 3.2). For measuring test case similarity, we compared several widely-used similarity metrics, including cosine similarity, Euclidean distance, and WMD. In addition, we investigated the impact of various GA initialization strategies on the overall performance of RTM by evaluating all three strategies introduced in Section 3.4.2.
@@
Overall, we observe that sentence-level embeddings generally outperform word-level embeddings in terms of FDR. Among the sentence-level embedding techniques, TF-IDF combined with cosine similarity achieves the best overall performance, with a peak FDR of $86.09\%$ under $PM3$ and Init Strategy 2. This suggests that even simple, sparse representations like TF-IDF can be highly effective for TSM, especially when the test cases consist of structured test steps that mainly differ in variable names and parameter values. When combined with appropriate preprocessing methods and initialization strategies, TF-IDF effectively captures the discriminative textual features in such test cases by leveraging word frequency. Similarly, Amazon Titan Text Embedding V2, LongT5, and USE also achieve competitive results, with the best Titan configuration reaching $81.59\%$ ($PM3$, Init Strategy 2, Euclidean distance), LongT5 achieving $78.50\%$ ($PM3$, Init Strategy 3, Euclidean distance), and USE achieving $78.05\%$ ($PM1$, Init Strategy 3, Euclidean distance).
@@
In terms of preprocessing, we observe that, for word-level embedding, $PM1$ and $PM2$ often yield better FDR results than $PM3$, indicating that preprocessing is required for using word-level embedding techniques. Word-level embedding is more sensitive to individual characters such as casing and punctuation, and thus proper preprocessing helps normalize these variations, enabling more accurate similarity computations. On the other hand, $PM1$ and $PM3$ tend to lead to higher FDR than using $PM2$ for sentence-level embeddings, meaning that less or no preprocessing can help these models better learn context information from test cases, without losing the key details that distinguish them.
@@ 5.2. Performance of RTM compared with baselines (RQ2)
Approach. To address RQ2, we compared the performance of RTM with the best configuration against baseline approaches in terms of FDR, varying the minimization budget from $10\%$ to $90\%$ in $10\%$ increments. According to RQ1, we selected RTM using $PM3$ as preprocessing strategy, TF-IDF as the embedding technique, Initialization Strategy 2 for GA initialization, and cosine similarity for the similarity measure as the best-performing configuration.
We evaluated RTM against Random Minimization, which also employs Initialization Strategy 2, and two versions of FAST-R (Cruciani et al., 2019), namely the fixed-budget version and the adequate version. For the fixed-budget version, we varied the minimization budget from $10\%$ to $90\%$, in increments of $10\%$. For the adequate scenario, we first executed FAST-R to determine its resulting minimization budget. This budget was then applied consistently to both RTM and the random minimization for experimental comparison. Notably, due to the redundancy in our dataset, as few as $7\%$ (i.e., $54$ test cases, one test case for each requirement) of the test cases are sufficient to achieve $100\%$ requirement coverage.
@@
Results. Figure 3 depicts the FDR for RTM, Random Minimization and the fixed-budget version of FAST-R, as well as the theoretical best possible FDR across minimization budgets from $10\%$ to $90\%$.
Overall, RTM consistently outperforms all other techniques across minimization budgets. The performance gap is particularly notable at mid-range budgets ($30\%$ to $60\%$). For instance, at a $40\%$ minimization budget, RTM achieves an FDR of $78.7\%$, whereas Random Minimization and FAST++ yield approximately $65.5\%$ and $56.5\%$, respectively. This demonstrates RTM’s superior ability to retain fault-revealing test cases, especially under mid-range budget constraints. At lower and higher minimization budgets, the FDR of all techniques tends to converge toward either 00 or $1$, thus making the difference in FDR between RTM and other techniques smaller. However, mid-range minimization budgets are more important in practice, as they offer more useful trade-offs between cost savings and fault detection.
As the minimization budget increases, the FDR of all techniques gradually improves and approaches the theoretical upper bound of $1.0$. Among the FAST-R variants, FAST-CS and FAST-all demonstrate the more competitive performance, while FAST++ and FAST-pw exhibit weaker results at lower budgets, indicating limited effectiveness in resource-constrained settings. Interestingly, Random Minimization consistently achieves higher FDR than FAST-R. This may be due to the distance metrics employed by FAST-R, which may not provide effective guidance for reducing similar test cases, whereas Random Minimization inherently preserves a greater diversity of test cases, leading to better FDR.
[…]
@@
RM with constraint 25.91 100.00 96.36 100.00
RM without constraint 19.86 39.26 95.41 98.52
RM: Random Minimization
In summary, compared with the baselines, RTM achieves the highest FDR across all minimization budgets, making it a more reliable choice for practical TSM with strong fault detection guarantees. In addition, RTM ensures full requirement coverage while enabling pre-defined minimization budgets, two features not simultaneously supported by FAST-R. These capabilities are particularly crucial in the context of requirement-based testing, where covering all requirements is often a strict necessity. Moreover, the ability to ensure fixed-budget minimization is essential in resource-constrained environments, where testing time or execution cost must be carefully managed.
Answering RQ2: RTM consistently outperforms all baselines in both FDR and requirement coverage under both the fixed-budget and adequate versions.
[…]
Figure 3. Comparison of FDR across minimization budgets for RTM, FAST-R, and Random Minimization, alongside the theoretical upper bound.
@@ 5.3. Impact of test suites redundancy (RQ3)
Approach. In RQ3, we investigate the impact of test suite redundancy levels on the performance (i.e.,FDR) of TSM techniques. Evaluating TSM techniques necessarily requires considering the test suite redundancy level, and we aim to shed light on this aspect. As described in Section 4.2, for each redundancy level (ranging from $4.5$ to $11.5$ in increments of $0.5$), we employ the ILP algorithm and GA to generate $10$ diverse test suites, each ensuring full coverage of both requirements and faults. We then apply RTM and the baseline methods to each of these test suites, and report the average FDR across the $10$ test suites for each redundancy level.
@@
Results. Figure 4 illustrates the FDR performance of RTM against baseline methods (FAST++, FAST-CS, FAST-all, FAST-pw, and Random Minimization with and without full requirement coverage constraint) across varying redundancy levels (ranging from $4.5$ to $11.5$ in increments of $0.5$) under seven distinct minimization budgets (from $30\%$ to $90\%$ in increments of $10\%$). Moreover, for each test suite, we calculate the theoretical best FDR for comparison purpose. Note that we cannot set minimization budgets below $30\%$ as some test suites are unable to achieve full requirement coverage under those constraints.
@@
We observe that RTM consistently achieves superior FDR compared to all baseline techniques across all redundancy levels and minimization budgets. Specifically, RTM more closely aligns with the theoretical optimum and consistently outperforms all other techniques. The difference in terms of FDR between RTM and baselines is particularly significant at mid-range minimization budgets ($30\%$$60\%$).
@@
As expected, the FDR of all techniques improves as the test suite redundancy level increases, indicating that the TSM techniques tend to perform better–achieving higher FDR–when applied to more redundant test suites. This trend, however, becomes less evident at higher budgets ($80\%$ and $90\%$), where the FDR of all techniques converge toward $1$. Random Minimization (with and without requirement coverage constraint) slightly outperforms other FAST-R variants but consistently falls short of RTM. We also observed that across all minimization budgets and redundancy levels, Random Minimization with the requirement constraint consistently outperformed the unconstrained variant in terms of FDR. This demonstrates that, for our dataset, explicitly ensuring that every requirement is covered not only preserves full requirement coverage but also significantly boosts the test suite’s ability to detect faults. In other words, prioritizing requirement coverage serves as an effective heuristic for enhancing fault detection under budgetary constraints.
@@
Moreover, the analysis reveals that redundancy level significantly influences FDR at lower minimization budgets (budgets below $60\%$), highlighting the importance of considering redundancy under tight budget constraints. Notably, higher redundancy combined with low minimization budgets tends to result in larger gaps between the FDR of all TSM techniques and the theoretical maximum. This discrepancy might be explained as under a very low minimization budget, the search space is severely restricted, making it challenging to assemble a diverse subset that both covers all requirements and maximizes fault coverage. Future research should account for redundancy levels when evaluating TSM techniques.
[…]
@@
Figure 4. Comparison of FDR Across Varying Redundancy Levels for RTM and Baseline Approaches Under Seven Minimization Budgets
@@
Answering RQ3: The test suite redundancy level has a strong effect, in general, on the achieved FDR of all TSM techniques. Further, RTM consistently outperforms all baseline techniques in terms of FDR across all redundancy levels and minimization budgets.

### v2

@@ 4.5. Evaluation Metrics
(15) $Coverage={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{|\mathcal{R^{\prime}}|}{|\mathcal{R}|}}$
where $\mathcal{R^{\prime}}$ represents the set of requirements covered by the minimized test suite $\mathcal{T^{\prime}}$, and $\mathcal{R}$ denotes the original set of requirements. The coverage value ranges from 0 to 1, where 0 indicates that the minimized test suite covers no requirements, while 1 signifies full coverage of all requirements.
@@ 4.6. Experiment Environment
All experiments were conducted on a computer provided by our industry partner, running Windows 10, equipped with an Intel Core i7-11850H CPU at 2.5 GHz and 32 GB of RAM. Due to data privacy considerations, we were restricted to performing our experiments solely on this laptop, limiting the text representation techniques available to us.
@@ 5.1. Performance of RTM under different configurations (RQ1)
Approach. To address RQ1, we systematically evaluated the performance of RTM under various configurations. Specifically, under a mid-range minimization budget ($50\%$), we ran RTM with three different preprocessing methods (as described in Section 3.1), and applied both word-level and sentence-level embedding techniques to transform test cases into vector representations (Section 3.2). To measure test case similarity, we compared several widely used metrics, including cosine similarity, Euclidean distance, and WMD. In addition, we investigated the impact of various GA initialization strategies on the overall performance of RTM by evaluating all three strategies introduced in Section 3.4.2.
@@
Overall, we observe that sentence-level embeddings generally outperform word-level embeddings in terms of FDR. Among the sentence-level embedding techniques, TF-IDF combined with cosine similarity achieves the best overall performance, with a peak FDR of $86.09\%$ under $PM3$ and Init Strategy 2. Similarly, Amazon Titan Text Embedding V2, LongT5, and USE also achieve competitive results, with the best Titan configuration reaching $81.59\%$ ($PM3$, Init Strategy 2, Euclidean distance), LongT5 achieving $78.50\%$ ($PM3$, Init Strategy 3, Euclidean distance), and USE achieving $78.05\%$ ($PM1$, Init Strategy 3, Euclidean distance).
Interestingly, despite being a simple, sparse representation, TF-IDF outperformed all other sentence-level embeddings. The characteristics of our industrial test cases can explain this. As shown in Figure 1, the test cases in our dataset follow structured and repetitive templates with little linguistic variation. The steps in the test cases differ only in variable names or parameter values, and TF-IDF precisely captures these differences through its word-frequency mechanism, down-weighting common words that appear in every test case and up-weighting rare words, including distinct variable names and parameter values. In contrast, other sentence-level embeddings, while capable of capturing deep semantic relationships, tend to smooth out subtle lexical differences that are crucial for distinguishing test cases in our dataset. This suggests that in template-based, domain-specific test descriptions, TF–IDF, which suppresses common terms and emphasizes rare tokens, often outperforms semantically rich embeddings that smooth over token-level distinctions.
@@
In terms of preprocessing, we observe that, for word-level embedding, $PM1$ and $PM2$ often yield better FDR results than $PM3$, indicating that preprocessing is required for using word-level embedding techniques. Word-level embeddings are more sensitive to individual characters, such as casing and punctuation, so proper preprocessing helps normalize these variations, enabling more accurate similarity computations. On the other hand, $PM1$ and $PM3$ tend to lead to higher FDR than using $PM2$ for sentence-level embeddings, meaning that less or no preprocessing can help these models better learn context information from test cases, without losing the key details that distinguish them.
@@
Note that we further conducted experiments under 25% and 75% minimization budgets. The results show that across these three minimization budgets, TF-IDF consistently yields the highest FDR. While the best preprocessing method, initialization strategy, and similarity measure could differ slightly across budgets, the corresponding FDR differences were within $0.025$ and do not affect the conclusions discussed above. This indicates that our configuration choice is not sensitive to the specific budget. Given that the $50\%$ minimization budget is a representative mid-range budget, we report the best configuration determined at $50\%$ and use it in subsequent comparisons. The results for other budgets are available in the replication package.
@@ 5.2. Performance of RTM compared with baselines (RQ2)
Approach. To address RQ2, we compared the performance of RTM with the best configuration against baseline approaches in terms of FDR, varying the minimization budget from $10\%$ to $90\%$ in $10\%$ increments. According to RQ1, we selected RTM using $PM3$ as preprocessing method, TF-IDF as the embedding technique, Initialization Strategy 2 for GA initialization, and cosine similarity for the similarity measure as the best-performing configuration.
We evaluated RTM against Random Minimization, which also employs Initialization Strategy 2, two versions of FAST-R (cruciani2019scalable)—the fixed-budget and adequate versions. For the fixed-budget version of FAST-R, we varied the minimization budget from $10\%$ to $90\%$, in increments of $10\%$. The fixed-budget version of FAST-R and RM-NoReq do not guarantee the 100% requirement coverage and are included in this RQ solely for FDR comparison. In the adequate scenario, we first ran FAST-R to determine the resulting minimization budget. This budget was then applied consistently to RTM and other baselines for experimental comparison. Notably, due to redundancy in our dataset, as few as $7\%$ (i.e., $54$ test cases, one test case per requirement) are sufficient to achieve $100\%$ requirement coverage.
@@
Results. Figure 3 depicts the FDR for RTM, Random Minimization, the fixed-budget version of FAST-R, as well as the theoretical best possible FDR across minimization budgets from $10\%$ to $90\%$.
We observe that RTM consistently outperforms Random Minimization and FAST-R approaches in terms of FDR across various minimization budgets. The performance gap is particularly notable at mid-range budgets ($30\%$ to $60\%$). For instance, at a $40\%$ minimization budget, RTM achieves an FDR of $78.7\%$, whereas RM-Req and FAST++ yield approximately $65.5\%$ and $56.5\%$, respectively. This demonstrates RTM’s superior ability to retain fault-revealing test cases, especially under mid-range budget constraints. At lower and higher minimization budgets, the FDR of all techniques tends to converge toward either 0 or $1$, thus making the difference in FDR between RTM and other techniques smaller. However, mid-range minimization budgets are more important in practice, as they offer more useful trade-offs between cost savings and fault detection.
As the minimization budget increases, the FDR of all techniques gradually improves and approaches the theoretical upper bound of $1.0$. Among the FAST-R variants, FAST-CS and FAST-all demonstrate better performance, while FAST++ and FAST-pw exhibit weaker results at lower budgets, indicating limited effectiveness in resource-constrained settings. Interestingly, Random Minimization generally achieves higher FDR than FAST-R. This may be due to the distance metrics employed by FAST-R, which may not provide effective guidance for reducing similar test cases, whereas Random Minimization inherently preserves a greater diversity of test cases, leading to better FDR. Moreover, the fixed-budget version of FAST-R does not guarantee full requirement coverage, making it unusable in our context.
For the adequate version of FAST-R, after reducing similarity among test cases while preserving requirement coverage of the minimized test suite, FAST++, FAST-CS, and FAST-pw ultimately retained only $7\%$ (i.e., selecting one test case under each requirement) of the test cases, while FAST-all retained $90\%$. We therefore compared RTM  and Random Minimization using $7\%$ and $90\%$ as the minimization budgets with the adequate version of FAST-R. The detailed results are presented in Table 2. Under the $7\%$ budget, RTM achieves a high FDR of $26.32\%$ while maintaining full requirement coverage ($100\%$), which is comparable to FAST++ ($27.27\%$) and FAST-CS ($25.95\%$), and outperforming FAST-pw in both FDR ($21.18\%$) and requirement coverage ($89.44\%$). RM-Req and RM-NoReq achieve FDR of $25.91\%$ and $19.86\%$, respectively.
Under the $90\%$ minimization budget, RTM achieves the highest FDR of $98.59\%$ among all techniques, maintaining full requirement coverage. In contrast, FAST-all shows slightly lower FDR ($92.59\%$) and slightly reduced coverage ($99.63\%$). RM-Req and RM-NoReq yield lower FDR of $96.36\%$ and $95.41\%$, respectively.
These results demonstrate that, RTM consistently delivers competitive FDR without compromising requirement coverage in the adequate scenario.
@@
RM-Req 25.91 100.00 96.36 100.00
RM-NoReq 19.86 39.26 95.41 98.52
In summary, compared with the baselines, RTM achieves the highest FDR on most minimization budgets, making it a more reliable choice for practical TSM with strong fault detection guarantees. In addition, RTM ensures full requirement coverage while enabling pre-defined minimization budgets, two features not simultaneously supported by FAST-R. These capabilities are particularly crucial in the context of requirement-based testing, where covering all requirements is often a strict necessity. Moreover, the ability to ensure fixed-budget minimization is essential in resource-constrained environments, where testing time or execution cost must be carefully managed.
Answering RQ2: RTM consistently outperforms all baselines in both FDR and requirement coverage on most minimization budgets.
Figure 3. Comparison of FDR across minimization budgets for RTM, FAST-R, and Random Minimization, alongside the theoretical upper bound.:
@@ 5.3. Impact of test suites redundancy (RQ3)
Approach. In RQ3, we investigate the impact of test suite redundancy levels on the performance (i.e.,FDR) of TSM techniques. Evaluating TSM techniques necessarily requires considering the test suite redundancy level, and we aim to shed light on this aspect. As described in Section 4.2, for each redundancy level (ranging from $4.5$ to $11.5$ in increments of $0.5$), we employ the ILP algorithm and GA to generate $10$ diverse test suites, each ensuring full coverage of both requirements and faults. We then apply RTM and the baseline techniques (Random Minimization and fixed-budget version of FAST-R) to each of these test suites, and report the average FDR across the $10$ test suites for each redundancy level. Note that the fixed-budget version of FAST-R and RM-NoReq do not guarantee 100% requirement coverage. We include them in this RQ solely for FDR comparison.
@@
Results. Figure 4 illustrates the FDR performance of RTM against baselines across varying redundancy levels (ranging from $4.5$ to $11.5$ in increments of $0.5$) under seven distinct minimization budgets (from $30\%$ to $90\%$ in increments of $10\%$). Moreover, for each test suite, we calculate the theoretical best FDR for comparison purposes. Note that we cannot set minimization budgets below $30\%$ because some test suites cannot achieve full requirement coverage under those constraints.
@@
We observe that RTM consistently achieves superior FDR compared to all baseline techniques across all redundancy levels and minimization budgets, while aligning more closely with the theoretical optimum. The difference in terms of FDR between RTM and baselines is particularly significant at mid-range minimization budgets (30%–60%).
@@
As expected, the FDR of all techniques improves as the test suite redundancy level increases, indicating that the TSM techniques tend to perform better–achieving higher FDR–when applied to more redundant test suites. This trend, however, becomes less evident at higher budgets ($80\%$ and $90\%$), where the FDR of all techniques converge toward $1$. Random Minimization (RM-Req and RM-NoReq) slightly outperforms other FAST-R variants but consistently falls short of RTM. We also observed that across all minimization budgets and redundancy levels, RM-Req consistently outperformed RM-NoReq in FDR. This demonstrates that, for our dataset, explicitly ensuring that every requirement is covered not only preserves full requirement coverage but also significantly boosts the test suite’s ability to detect faults. In other words, prioritizing requirement coverage is an effective heuristic for improving fault detection under budgetary constraints.
@@
Moreover, the analysis reveals that redundancy level significantly influences FDR at lower minimization budgets (below $60\%$), highlighting the importance of considering redundancy under tight budget constraints. Notably, higher redundancy combined with low minimization budgets tends to result in larger gaps between the FDR of all TSM techniques and the theoretical maximum. This discrepancy might be explained as follows: under a very low minimization budget, the search space is severely restricted, making it challenging to assemble a diverse subset that both covers all requirements and maximizes fault coverage. Future research should account for redundancy levels when evaluating TSM techniques.
@@
Figure 4. Comparison of FDR Across Varying Redundancy Levels for RTM and Baseline Approaches Under Seven Minimization Budgets:
@@
Answering RQ3: The test suite redundancy level generally has a strong effect on the achieved FDR of all TSM techniques. Further, RTM consistently outperforms all baseline techniques in terms of FDR across all redundancy levels and minimization budgets.
5.4. Scalability of RTM compared with baselines (RQ4)
Approach. To address RQ4, we compared the performance of RTM with the best configuration against baselines in terms of scalability, under $50\%$ minimization budget. As described in Section 4.2, we generate test suites under $15$ redundancy levels with various test suite sizes, resulting in a total of $150$ test suites. To investigate how RTM scales with test suite size compared to baselines, we collect the runtime, including preparation time (time for embeddings generation and similarity calculation), search time (time for searching the optimal subsets), and total minimization time (the summation of preparation time and search time), across $150$ test suites with different test suite sizes. Moreover, to quantify and compare the relationships between runtime and test suite size across different approaches, we fit regression models for preparation time, search time, and total minimization time as functions of test suite size.
Results. Figure 5 depicts how the preparation time, search time, and total minimization time of RTM and baseline approaches scale with the test suite sizes (number of test cases) under a $50\%$ minimization budget. We observe that the observations in Figure 5 approximately follow a linear relationship. Therefore, for each approach, a linear regression model is fitted to the preparation time, search time, and total minimization time as functions of test suite size. The equation of the linear regression model is as follows:
(16) Time=an+cTime=a*n+c
where $n$ denotes the number of test cases, $a$ is the regression coefficient of the linear term, $c$ denotes the intercept of the model.
The results of the linear regression model confirm that the total minimization time, preparation time, as well as search time and the number of test cases, for RTM and baseline approaches, follow a linear relationship with the coefficient on the linear term statistically significant with $\alpha=0.01$. The $R^{2}$ values for all the regression models are above $93\%$, indicating a high percentage of variation in runtime can be explained by the number of test cases.
We observe that, although being slower than all other baselines, RTM minimizes test cases in seconds–$1.32$s to $4.44$s in terms of total minimization time–for test suite size from $187$ to $616$, which is computationally efficient. Moreover, the fact that the total minimization time of RTM follows a linear relationship with the test suite size indicates that RTM is scalable for larger test suites. According to the fitted linear regression model, when the test suite size $n$ is 200,000200{,}000, the predicted total minimization time of RTM is $6.217686\times 10^{-3}\,n+2.523669\approx 1246.06$ seconds, whereas FAST-CS (the fastest FAST-R approach) is $1.195291\times 10^{-4}\,n+0.051979\approx 23.96$ seconds and RM-Req is $9.062634\times 10^{-7}*n+0.000472\approx 0.18$ seconds. Thus, the time differences are $20.37$ mins and $20.76$ mins when compared to FAST-CS and RM-Req, respectively. Given the fact that, unlike test case selection and prioritization, test suite minimization only performs occasionally, such as major releases with many new added test cases and redundant old test cases (DBLP:journals/tse/PanGB24; pan2023atm), such differences are practically acceptable.
Moreover, the preparation time of RTM ranges from $0.10$s to $0.67$s, which is negligible in our context. We also observe that RTM is much faster than FAST-all in terms of preparation time. According to the regression model results, when the test suite size is 200,000200{,}000, the predicted absolute preparation time difference between RTM and FAST-CS is $3.30$ min, which is of no practical consequence for TSM. As for the search time, RTM took $1.21$s to $3.85$s across $150$ test suites. When the test suite size is 200,000200{,}000, the predicted absolute search time difference of RTM  when compared to FAST-CS is $17.07$ mins.
In summary, RTM runs within $5$ seconds across all evaluated test suites. Although being slower than the baselines, RTM scales approximately linearly with test suite sizes, and even for substantially larger suites, the absolute runtime gaps remain acceptable in practice—they are not expected to have practical consequences for TSM. Given that RTM consistently achieves much higher FDR across various datasets and minimization budgets than the baselines while preserving full requirement coverage, it is therefore a better option in many practical contexts.
Answering RQ4: RTM is scalable in terms of runtime and, on the evaluated datasets, exhibits small absolute time differences (i.e., of no practical consequences) compared to the baselines while achieving much higher FDR performance with full requirement coverage preserved, thus making it a better choice in practice.
Figure 5. Scatter plots of the number of test cases and preparation time, search time and total minimization time (in sec) for RTM and baselines under 50% minimization budget: